# The Origin of Quantum Mechanical Statistics: Some Insights from the Research on Human Language

### v1

@@ 5 The appearance of Bose–Einstein statistics in language
We do not expound the theoretical scheme in detail here and limit ourselves to summarize the main results, for the sake of brevity (for the full development of these ideas, see, e.g., (Aerts & Beltran,, 2020; Beltran,, 2021; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ) and references therein). We instead aim, in this and the following section, to explain why Bose–Einstein statistics arises in human language and to reflect about the specific lack of statistical independence that Einstein identified in the behaviour of a gas of photons and how in cognitive-linguistics domains there could be a connection with a ‘mechanism in the dynamics of meaning’ that occurs whenever words, and their respective concepts, are combined to produce a text of human language.
We preliminarily stress that we consider words as labels for concepts in our theoretical scheme, that is, they are the concepts that give the words they refer to their meaning. More precisely, we consider any concept as an abstract entity of meaning that is at a given time in a given state, and the latter is what captures the meaning content of the concept. Individual concepts can then be composed, or combined, to form more complex entities of meaning, e.g., conceptual combinations. Proceeding in this way, a written text is an entity of meaning obtained by combining the concepts that correspond to all the words that appear in the text. We have demonstrated in several articles that the meaning of a written text produced by human language relates to the meaning of the words that appear in the text in a complex way which deviates from the prescriptions of classical logical semantics (see, e.g., (Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ; Philtransa,, 2023)).
@@
Why shall we expect Bose–Einstein statistics to hold in conceptual combinations? To understand this point, let us firstly consider the simple concept combination Eleven Animals. It is clear that, at a conceptual level, each one of the eleven animals is completely ‘identical with’ and ‘indistinguishable from’ each other of the eleven animals. On the other side, it is also clear that, in the case of ‘eleven physical animals’, there are always differences between each of them because, as ‘objects’ present in the physical world, they have an individuality and, as individuals with spatially localized physical bodies, they can be distinguished from each other. In the latter case, we expect Maxwell–Boltzmann statistics to apply. In the former case, we instead expect a non-classical statistics to apply as a consequence of the fact that, due to their conceptual nature, the eleven animals are intrinsically indistinguishable. We have proved in (Aerts & Sozzo & Veloz,, 2015) that, indeed, Bose–Einstein statistics better models the concept combination Eleven Animals as compared to Maxwell–Boltzmann.
@@
Relying on the above insights, we have analysed in detail a story-telling text, namely, the Winnie the Pooh story entitled “In Which Piglet Meets a Haffalump” (Milne,, 1926) and we have calculated the frequency of appearance of all the words in the text. Next, we have interpreted these frequencies of appearance as the cognitive counterpart of single-entity energy levels, hence micro-states, of identical physical entities.
We have found that these energy levels significantly differ from the Maxwell–Boltzmann distribution in Equation (5), whereas they perfectly distribute according to the Bose–Einstein distribution in Equation (4). This has allowed us to conclude that a text of human language can be considered in terms of its meaning content as an ideal gas of identical and indistinguishable conceptual entities, which we have called ‘cognitons’ (see Section 1), in complete analogy with the case of an ideal gas of bosons (Aerts & Beltran,, 2020; Beltran,, 2021; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ). Furthermore, we have provided arguments to support the hypothesis that the gas of cognitons is in an overall state that is close to a Bose–Einstein condensate (Aerts & Beltran,, 2020). The analysis has been implemented on various texts, including short and long stories, e.g., novels, and we always found the results that we summarize in the following.
@@
In our theoretical scheme, the numbers of appearance $N(E_{i})$, $i=0,1,\ldots,n$, in Equations (6) and (7) can be used to determine the pairs of constants $(A,B)$ and $(C,D)$ in Equation (4) and (5), respectively, subject to the constants $N=2,655$ and $E=242,891$, thus checking whether the Bose–Einstein distribution or the Maxwell–Boltzmann distribution fit the data. This task has effectively been accomplished in (Aerts & Beltran,, 2020; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ), comparing the ensuing Bose–Einstein and Maxwell–Boltzmann distributions with empirical data. Figures 1 and 2 summarise the comparison.
@@
Figure 1: In (a) we represent the number of appearances $N(E_{i})$ of words in the Winnie the Pooh story “In Which Piglet Meets a Haffalump” (Milne,, 1926), ranked from lowest energy level, corresponding to the most often appearing word, to highest energy level, corresponding to the least often appearing word. The blue graph (Series 1) represents the data, i.e. the collected numbers of appearances from the story, the red graph (Series 2) is a Bose–Einstein distribution model for these numbers of appearances, and the green graph (Series 3) is a Maxwell–Boltzmann distribution model. In (b) we represent the $\log/\log$ graphs of the numbers of appearances and their Bose–Einstein and Maxwell–Boltzmann models. The red and blue graphs coincide almost completely in both plots, whereas the green graph does not coincide at all with the blue graph of the data. This shows that the Bose–Einstein statistical distribution is a good model for the numbers of appearances, while the Maxwell–Boltzmann distribution is not.
Figure 2: The figure reports the energy distribution of the Winnie the Pooh story “In Which Piglet Meets a Haffalump” (Milne,, 1926). The blue graph (Series 1) represents the energy $E_{i}N(E_{i})$ radiated by the story per energy level $E_{i}=i$, the red graph (Series 2) represents that same energy per energy level as modelled by the Bose–Einstein distribution, and same for the green graph (Series 3), which follows the Maxwell–Boltzmann distribution.
The results of the comparison significantly reveal that ‘Bose–Einstein statistics is in remarkably good fit with empirical data’, whereas a ‘large deviation is observed in the data from Maxwell–Boltzmann statistics’. The conclusion is that, once a notion of energy is adequately introduced and quantified from empirical data, a collection of words in a written text produced by human language behaves as a suitable gas of identical bosons. As a matter of fact, we have introduced the term ‘cogniton’ as the fundamental quantum of cognition. Equivalently, each word $w_{i}$ with energy $E_{i}$ corresponds to a micro-state of a cogniton with energy $E_{i}$. The overall text then behaves as a ‘gas of cognitons’ whose energies are distributed according to Bose–Einstein statistics. [^7] It is worth to mention, at this stage, that we have also provided a theoretical foundation of ‘Zipf’s law’ in human language (Zipf,, 1935, 1949), whose theoretical origin is not understood yet and whose nature is therefore considered to be merely empirical. On the contrary, our results in (Aerts & Beltran,, 2020; Aerts & Beltran, 2022a, ; Aerts & Beltran, 2022b, ) derive it on the basis of the presence of a Bose–Einstein behaviour in the energy levels of words in a text.
@@
It would of course be possible to test the specific example we have considered using the calibrated method in psychology, with a suitably chosen Likert scale, then make an experimental estimate of the statistics with which children would choose between two kittens, two puppies or a kitten and a puppy. This specific experiment was however not carried out by us, but we examined very similar experimental situations, which also included a choice between cats and dogs, being able to show that the Bose–Einstein statistics systematically constitutes a better model than the Maxwell–Boltzmann statistics (Aerts & Sozzo & Veloz,, 2015; Aerts & Beltran,, 2020).
The phenomenon we have called ‘contextual updating’ contains a possible explanation for this statistical behaviour. In fact, the conceptual combinations ‘two kittens’, ‘two puppies’, and ‘a kitten and a puppy’, each behave differently with respect to the entire context of meaning that characterizes a child’s choosing, and in this sense they acquire their own individual probabilities of occurrence. Does this mean that the probabilities of these choices are necessarily equal and that a Bose–Einstein statistics follows? Not necessarily, contextual updating will only cause them to be associated with non-decomposable probabilities, thus failing the well-known property, typical of the Maxwell–Boltzmann statistics, that the combination ‘one kitten and one puppy’ possesses twice the frequency of ‘two kittens’ and ‘two puppies’. In other words, the statistical independence of the choices ‘kitten’ and ‘puppy’ is lost as a consequence of a mechanism, that of contextual updating, that we identified as lying also at the origin of the violation of Bell’s inequalities (Ijtp,, 2023).
Although we will not elaborate on this in this article, we would like to point out that the fact that Bose–Einstein statistics prescribes the same probabilities of occurrence for the different conceptual combinations is perhaps due to a statistical averaging effect, similar to the ‘universal average’ that we introduced and studied in relation to Born probabilities (Aerts, Sassoli,, 2014, 2015). Indeed, if the contextual updating mechanism underlies the loss of statistical independence, we can imagine that each of the possible combinations could elicit a multitude of different meanings. In our example, a child choosing may be influenced by the fact that adult cats and dogs are often not friends with each other, making the choice ‘a kitten and a puppy’ less attractive than the choice ‘two kittens’ or ‘two puppies’. However, there are also contexts of meaning that make this choice more attractive, for example, when a child wants to play both with a kitten and a puppy. This multitude of meaning contexts correspond to ‘different ways of choosing’ the available conceptual combinations, and each way of choosing is associated with a different statistics. When an average over them is considered, this is expected to lead to a Bose–Einstein type of quantum statistics (Aerts, Sassoli,, 2014, 2015).
@@ 6 Randomness and temperature
We intend to explore in this section the role that ‘meaning’ might play in this lack of statistical independence. Let us support our analysis with an example and consider again the Winnie the Pooh story “In which Piglet meets a Haffalump” which we have discussed in Section 5. The word Piglet occurs $47$ times in the story, while the word First occurs $7$ times. Let us now suppose that one would have asked the author to write some additional paragraphs to the story text. One realizes at once that the probability that the word Piglet would occur in these additional paragraphs is higher than the probability that the word First appears in them. The reason is that the story text carries an overall meaning, hence the words that have more affinity with such overall meaning have a higher chance of appearing when the story is continued to be written. Equivalently, one can say that ‘meaning is a force that makes the same words attract each other to clump together as a result of this meaning force’ (Aerts & Beltran, 2022b, ). Hence, meaning could be what plays the conceptual counterpart of the force that Einstein believed to cause photons to clump into the same state. Each word that is added to the text is connected with all the words already existing in the text if the latter is telling a story that possesses coherence in terms of meaning, as is usually the case for stories told by human beings.
@@
We repeated this analysis of introducing an element of randomness, consisting of spreading the numbers of appearance of the words originally used in a story over its placeholders, for the other stories we studied in (Aerts & Beltran,, 2020), entitled “The Magic Shop” and “Gulliver’s Travels”. And we obtained even more data for this analysis of heat and temperature by implementing this randomness as well on several other stories of novels we studied using corpora of the Italian language (Italiancorpora,, 2024). The result was the same, namely, the total energy and temperature of each of the stories increased in a way that is similar to that of the Winnie the Pooh story. We do not give here details of our analysis of the thermodynamic temperature of stories in human language and refer to (Italiancorpora,, 2024; Temperature,, 2024) for the complete study, but draw the following conclusion as to its importance for the subject of our article. Even when a specific part of the meaning contained in a story is taken away and the words are placed at random in a new text, a Bose–Einstein statistical model still exists for it, which means that even these more random collections of words carry the phenomenon of condensation, that is, the clumping together of the same words.
@@
However, it should be added that we have introduced randomness into the original story by using the already existing fabric of words that form it, and this can be interpreted as implementing a certain type of locality. In fact, placeholders for words in the original story can be regarded as precursors of a ‘locality claimed by words’. Note in this regard that when we index the placeholders, those words that would otherwise be absolutely indistinguishable from each other actually become distinguishable (for example, to the fourth time that ‘Piglet’ appears in the story of Winnie the Pooh, we can assign index 4). In this sense, the specific situation in which only the words of the original story are subjected to randomisation may be similar to the situation of a gas that has been entirely localised in a container in a given physics laboratory. Indeed, molecules of the same gas located in another container, located in another laboratory, will play no role in determining its temperature. If we consider experiments in which a Bose–Einstein condensate is made from a specific gas, and let the temperature slowly rise again, we will go through a phase in which quantum coherence is indeed still partially present, just as in the randomisation we introduced ‘meaning’ is still partially present. In this regard, it is interesting to ask what phase a star like our sun is currently in (Temperature,, 2024).
@@ 7 The possible role of meaning
Let us first make it clear that we do not consider here all aspects of how meaning plays a role in the dynamics that takes place and in the structure that is present in human language, also because several of these aspects are currently the subject of ongoing research. However, already at this stage, we can draw from our analysis in Sections 5 and 6 some interesting explanatory hypothesis and insights. There is a specific mechanism that we identified in one of our previous publications (Ijtp,, 2023), which we called ‘contextual updating’, that incorporates an important part of what meaning does.
More precisely, we have proved in several articles that conceptual meaning can be represented by entanglement in the Hilbert space formalism of QM (see, e.g., (Aerts, Gabora, Sozzo et al,, 2013; Aerts & Sozzo,, 2016; Aerts, Sassoli, Sozzo,, 2016; Pisano & Sozzo,, 2020; Aerts et al.,, 2021; Aerts & Beltran, 2022a, ; Philtransa,, 2023; Ijtp,, 2023)). Furthermore, we have recently suggested that, in both language and physics domains, entanglement can be considered as a ‘phenomenon of contextual updating’ (Ijtp,, 2023; Philtransa,, 2023). According to this view, in both language and physics domains, the lack of statistical independence would be produced by a phenomenon of entanglement through contextual updating.
To explain what we mean by this mechanism of contextual updating, let us consider the combination of two words in a text. Each word carries meaning but, whenever the two words are combined, their combination also carries its proper meaning, which is not the trivial combination of the meanings associated with the two individual words as prescribed by a classical logical semantics. The new emergent meaning of the combined word arises in a complex contextual way, in which the whole of the context relevant to the story plays a fundamental role. Thus, each time a word is added to a text, a mechanism of updating influenced by the meaning of the whole context occurs, and this updating continues to take place until the end of the story that contains all the words. We proved that this mechanism of contextual updating through which meaning is attributed has to be carried by an entangled state (Ijtp,, 2023). For those familiar with the Hilbert space formalism of QM, and the ensuing tensor product for the description of composite quantum entities, they will recognize such a contextual updating mechanism in the mathematical procedure of describing multiple composite quantum entities. Indeed, whenever a Hilbert space of an individual quantum entity is coupled to other state spaces via the tensor product, new states form as a consequence of the superposition principle, which always contains a majority of entangled states. It is these superposition states that accomplish the contextual updating in the mathematical formalism.
@@
This mechanism of contextual updating is very characteristic of how ‘meaning attaches and interweaves itself with’ a collection of words. In general, the writer’s intention is to make the text as clear as possible, which in our thermodynamic formalism translates into a ‘pure state’, with von Neumann entropy equal to zero. Whether this pure state is also pure for the reader and/or listener of the text is an interesting question in itself about which, however, we will not elaborate in this article (Temperature,, 2024).
@@
Hence, in general, meaning, as used by people in human language, attaches and interweaves with states that are close to this Bose–Einstein condensate but usually not equal to it. It is an intriguing, but not yet experimentally solved question, whether things happen similarly in the micro-world in relation to ‘quantum coherence’, which would play an equivalent role to the notion of meaning in the human cognitive domain (Temperature,, 2024). In any case, we believe that the analysis we have put forward in the present article comes closer to an explanation than the explanation produced by the symmetrization postulate and the ensuing complete indistinguishability of identical quantum physical entities.
@@
To conclude, we have observed that the classical statistical interpretation of thermodynamics rests on the assumption that physical entities behave in a random and independent way at a microscopic level, which entails maximization of entropy (Huang,, 1987). Based on our analysis of human language, we suggested that it is more appropriate instead to consider that, also in the case of physics, there is no situation where entities are truly random and independent, as (conceptual or physical) entities always bundle (entangle) together due to the presence of Bose–Einstein statistics, with the aim to collaborate and reduce the overall uncertainty, hence entropy, and produce globally a pure state (Ijtp,, 2023; Philtransa,, 2023).
@@
The results above clearly indicate that a theoretical direction towards the development of a quantum-type thermodynamics that is alternative to existing formal approaches to ‘quantum thermodynamics for physics’, none of which is satisfying (Gemmer et al.,, 2009; Mahler,, 2015; Gogolin & Eisert,, 2016), is a valuable aim to pursue. A possible direction for such an alternative theoretical approach rests on the notions of energy and entropy as defined in cognitive-linguistics domains, and we refer to (Philtransa,, 2023) for more details about the development of a thermodynamics of human cognition and human culture.

### v2

@@ 5 The appearance of Bose–Einstein statistics in language
We do not expound the theoretical scheme in detail here and limit ourselves to summarize the main results, for the sake of brevity (for the full development of these ideas, see, e.g., [38, 39, 40, 41] and references therein). We instead aim, in this and the following section, to explain why Bose–Einstein statistics arises in human language and to reflect about the specific lack of statistical independence that Einstein identified in the behaviour of a gas of photons and how in cognitive-linguistics domains there could be a connection with a ‘mechanism in the dynamics of meaning’ that occurs whenever words, and their respective concepts, are combined to produce a text of human language.
We preliminarily stress that we consider words as labels for concepts in our theoretical scheme, that is, they are the concepts that give the words they refer to their meaning. More precisely, we consider any concept as an abstract entity of meaning that is at a given time in a given state, and the latter is what captures the meaning content of the concept. Individual concepts can then be composed, or combined, to form more complex entities of meaning, e.g., conceptual combinations. Proceeding in this way, a written text is an entity of meaning obtained by combining the concepts that correspond to all the words that appear in the text. We have demonstrated in several articles that the meaning of a written text produced by human language relates to the meaning of the words that appear in the text in a complex way which deviates from the prescriptions of classical logical semantics (see, e.g., [40, 41, 43]).
@@
Why shall we expect Bose–Einstein statistics to hold in conceptual combinations? To understand this point, let us firstly consider the simple concept combination Eleven Animals. It is clear that, at a conceptual level, each one of the eleven animals is completely ‘identical with’ and ‘indistinguishable from’ each other of the eleven animals. On the other side, it is also clear that, in the case of ‘eleven physical animals’, there are always differences between each of them because, as ‘objects’ present in the physical world, they have an individuality and, as individuals with spatially localized physical bodies, they can be distinguished from each other. In the latter case, we expect Maxwell–Boltzmann statistics to apply. In the former case, we instead expect a non-classical statistics to apply as a consequence of the fact that, due to their conceptual nature, the eleven animals are intrinsically indistinguishable. We have proved in [70] that, indeed, Bose–Einstein statistics better models the concept combination Eleven Animals as compared to Maxwell–Boltzmann.
@@
Relying on the above insights, we have analysed in detail a story-telling text, namely, the Winnie the Pooh story entitled “In Which Piglet Meets a Haffalump” [71] and we have calculated the frequency of appearance of all the words in the text. Next, we have interpreted these frequencies of appearance as the cognitive counterpart of single-entity energy levels, hence micro-states, of identical physical entities.
We have found that these energy levels significantly differ from the Maxwell–Boltzmann distribution in Equation (5), whereas they perfectly distribute according to the Bose–Einstein distribution in Equation (4). This has allowed us to conclude that a text of human language can be considered in terms of its meaning content as an ideal gas of identical and indistinguishable conceptual entities, which we have called ‘cognitons’ (see Section 1), in complete analogy with the case of an ideal gas of bosons [38, 39, 40, 41]. Furthermore, we have provided arguments to support the hypothesis that the gas of cognitons is in an overall state that is close to a Bose–Einstein condensate [38]. The analysis has been implemented on various texts, including short and long stories, e.g., novels, and we always found the results that we summarize in the following.
@@
In our theoretical scheme, the numbers of appearance $N(E_{i})$, $i=0,1,\ldots,n$, in Equations (6) and (7) can be used to determine the pairs of constants $(A,B)$ and $(C,D)$ in Equation (4) and (5), respectively, subject to the constants $N=2,655$ and $E=242,891$, thus checking whether the Bose–Einstein distribution or the Maxwell–Boltzmann distribution fit the data. This task has effectively been accomplished in [38, 40, 41], comparing the ensuing Bose–Einstein and Maxwell–Boltzmann distributions with empirical data. Figures 1 and 2 summarise the comparison.
@@
Figure 1: In (a) we represent the number of appearances $N(E_{i})$ of words in the Winnie the Pooh story “In Which Piglet Meets a Haffalump” [71], ranked from lowest energy level, corresponding to the most often appearing word, to highest energy level, corresponding to the least often appearing word. The blue graph (Series 1) represents the data, i.e. the collected numbers of appearances from the story, the red graph (Series 2) is a Bose–Einstein distribution model for these numbers of appearances, and the green graph (Series 3) is a Maxwell–Boltzmann distribution model. In (b) we represent the $\log/\log$ graphs of the numbers of appearances and their Bose–Einstein and Maxwell–Boltzmann models. The red and blue graphs coincide almost completely in both plots, whereas the green graph does not coincide at all with the blue graph of the data. This shows that the Bose–Einstein statistical distribution is a good model for the numbers of appearances, while the Maxwell–Boltzmann distribution is not.
Figure 2: The figure reports the energy distribution of the Winnie the Pooh story “In Which Piglet Meets a Haffalump” [71]. The blue graph (Series 1) represents the energy $E_{i}N(E_{i})$ radiated by the story per energy level $E_{i}=i$, the red graph (Series 2) represents that same energy per energy level as modelled by the Bose–Einstein distribution, and same for the green graph (Series 3), which follows the Maxwell–Boltzmann distribution.
The results of the comparison significantly reveal that ‘Bose–Einstein statistics is in remarkably good fit with empirical data’, whereas a ‘large deviation is observed in the data from Maxwell–Boltzmann statistics’. The conclusion is that, once a notion of energy is adequately introduced and quantified from empirical data, a collection of words in a written text produced by human language behaves as a suitable gas of identical bosons. As a matter of fact, we have introduced the term ‘cogniton’ as the fundamental quantum of cognition. Equivalently, each word $w_{i}$ with energy $E_{i}$ corresponds to a micro-state of a cogniton with energy $E_{i}$. The overall text then behaves as a ‘gas of cognitons’ whose energies are distributed according to Bose–Einstein statistics. [^7] It is worth to mention, at this stage, that we have also provided a theoretical foundation of ‘Zipf’s law’ in human language [72, 73], whose theoretical origin is not understood yet and whose nature is therefore considered to be merely empirical. On the contrary, our results in [38, 40, 41] derive it on the basis of the presence of a Bose–Einstein behaviour in the energy levels of words in a text.
@@
It would of course be possible to test the specific example we have considered using the calibrated method in psychology, with a suitably chosen Likert scale, then make an experimental estimate of the statistics with which children would choose between two kittens, two puppies or a kitten and a puppy. This specific experiment was however not carried out by us, but we examined very similar experimental situations, which also included a choice between cats and dogs, being able to show that the Bose–Einstein statistics systematically constitutes a better model than the Maxwell–Boltzmann statistics [70, 38].
The phenomenon we have called ‘contextual updating’ contains a possible explanation for this statistical behaviour. In fact, the conceptual combinations ‘two kittens’, ‘two puppies’, and ‘a kitten and a puppy’, each behave differently with respect to the entire context of meaning that characterizes a child’s choosing, and in this sense they acquire their own individual probabilities of occurrence. Does this mean that the probabilities of these choices are necessarily equal and that a Bose–Einstein statistics follows? Not necessarily, contextual updating will only cause them to be associated with non-decomposable probabilities, thus failing the well-known property, typical of the Maxwell–Boltzmann statistics, that the combination ‘one kitten and one puppy’ possesses twice the frequency of ‘two kittens’ and ‘two puppies’. In other words, the statistical independence of the choices ‘kitten’ and ‘puppy’ is lost as a consequence of a mechanism, that of contextual updating, that we identified as lying also at the origin of the violation of Bell’s inequalities [42].
Although we will not elaborate on this in this article, we would like to point out that the fact that Bose–Einstein statistics prescribes the same probabilities of occurrence for the different conceptual combinations is perhaps due to a statistical averaging effect, similar to the ‘universal average’ that we introduced and studied in relation to Born probabilities [74, 75]. Indeed, if the contextual updating mechanism underlies the loss of statistical independence, we can imagine that each of the possible combinations could elicit a multitude of different meanings. In our example, a child choosing may be influenced by the fact that adult cats and dogs are often not friends with each other, making the choice ‘a kitten and a puppy’ less attractive than the choice ‘two kittens’ or ‘two puppies’. However, there are also contexts of meaning that make this choice more attractive, for example, when a child wants to play both with a kitten and a puppy. This multitude of meaning contexts correspond to ‘different ways of choosing’ the available conceptual combinations, and each way of choosing is associated with a different statistics. When an average over them is considered, this is expected to lead to a Bose–Einstein type of quantum statistics [74, 75].
@@ 6 Randomness and temperature
We intend to explore in this section the role that ‘meaning’ might play in this lack of statistical independence. Let us support our analysis with an example and consider again the Winnie the Pooh story “In which Piglet meets a Haffalump” which we have discussed in Section 5. The word Piglet occurs $47$ times in the story, while the word First occurs $7$ times. Let us now suppose that one would have asked the author to write some additional paragraphs to the story text. One realizes at once that the probability that the word Piglet would occur in these additional paragraphs is higher than the probability that the word First appears in them. The reason is that the story text carries an overall meaning, hence the words that have more affinity with such overall meaning have a higher chance of appearing when the story is continued to be written. Equivalently, one can say that ‘meaning is a force that makes the same words attract each other to clump together as a result of this meaning force’ [41]. Hence, meaning could be what plays the conceptual counterpart of the force that Einstein believed to cause photons to clump into the same state. Each word that is added to the text is connected with all the words already existing in the text if the latter is telling a story that possesses coherence in terms of meaning, as is usually the case for stories told by human beings.
@@
We repeated this analysis of introducing an element of randomness, consisting of spreading the numbers of appearance of the words originally used in a story over its placeholders, for the other stories we studied in [38], entitled “The Magic Shop” and “Gulliver’s Travels”. And we obtained even more data for this analysis of heat and temperature by implementing this randomness as well on several other stories of novels we studied using corpora of the Italian language [76]. The result was the same, namely, the total energy and temperature of each of the stories increased in a way that is similar to that of the Winnie the Pooh story. We do not give here details of our analysis of the thermodynamic temperature of stories in human language and refer to [76, 47] for the complete study, but draw the following conclusion as to its importance for the subject of our article. [^8] It is remarkable that, despite the specific linguistic differences between English and Italian languages, we have identified in the study of Italian novels a statistical behaviour that is completely analogous to what we have found in English texts, with respect to both Bose–Einstein distribution and appearance of effects due to temperature in the presence of randomization. This preliminary analysis shows that there are general meaning-related patterns in human language that are ‘species-specific’. Hence, we expect that similar patterns can also be identified if different languages are analysed. Even when a specific part of the meaning contained in a story is taken away and the words are placed at random in a new text, a Bose–Einstein statistical model still exists for it, which means that even these more random collections of words carry the phenomenon of condensation, that is, the clumping together of the same words.
@@
However, it should be added that we have introduced randomness into the original story by using the already existing fabric of words that form it, and this can be interpreted as implementing a certain type of locality. In fact, placeholders for words in the original story can be regarded as precursors of a ‘locality claimed by words’. Note in this regard that when we index the placeholders, those words that would otherwise be absolutely indistinguishable from each other actually become distinguishable (for example, to the fourth time that ‘Piglet’ appears in the story of Winnie the Pooh, we can assign index 4). In this sense, the specific situation in which only the words of the original story are subjected to randomisation may be similar to the situation of a gas that has been entirely localised in a container in a given physics laboratory. Indeed, molecules of the same gas located in another container, located in another laboratory, will play no role in determining its temperature. If we consider experiments in which a Bose–Einstein condensate is made from a specific gas, and let the temperature slowly rise again, we will go through a phase in which quantum coherence is indeed still partially present, just as in the randomisation we introduced ‘meaning’ is still partially present. In this regard, it is interesting to ask what phase a star like our sun is currently in [47].
@@ 7 The possible role of meaning
Let us first make it clear that we do not consider here all aspects of how meaning plays a role in the dynamics that takes place and in the structure that is present in human language, also because several of these aspects are currently the subject of ongoing research. However, already at this stage, we can draw from our analysis in Sections 5 and 6 some interesting explanatory hypothesis and insights. There is a specific mechanism that we identified in one of our previous publications [42], which we called ‘contextual updating’, that incorporates an important part of what meaning does.
More precisely, we have proved in several articles that conceptual meaning can be represented by entanglement in the Hilbert space formalism of QM (see, e.g., [28, 17, 18, 19, 20, 40, 43, 42]). Furthermore, we have recently suggested that, in both language and physics domains, entanglement can be considered as a ‘phenomenon of contextual updating’ [42, 43]. According to this view, in both language and physics domains, the lack of statistical independence would be produced by a phenomenon of entanglement through contextual updating.
To explain what we mean by this mechanism of contextual updating, let us consider the combination of two words in a text. Each word carries meaning but, whenever the two words are combined, their combination also carries its proper meaning, which is not the trivial combination of the meanings associated with the two individual words as prescribed by a classical logical semantics. The new emergent meaning of the combined word arises in a complex contextual way, in which the whole of the context relevant to the story plays a fundamental role. Thus, each time a word is added to a text, a mechanism of updating influenced by the meaning of the whole context occurs, and this updating continues to take place until the end of the story that contains all the words. We proved that this mechanism of contextual updating through which meaning is attributed has to be carried by an entangled state [42]. For those familiar with the Hilbert space formalism of QM, and the ensuing tensor product for the description of composite quantum entities, they will recognize such a contextual updating mechanism in the mathematical procedure of describing multiple composite quantum entities. Indeed, whenever a Hilbert space of an individual quantum entity is coupled to other state spaces via the tensor product, new states form as a consequence of the superposition principle, which always contains a majority of entangled states. It is these superposition states that accomplish the contextual updating in the mathematical formalism. [^9] For a detailed investigation of the problems of separability and entanglement in both quantum physics and human language within axiomatic foundations of quantum mechanics, see also [77].
@@
This mechanism of contextual updating is very characteristic of how ‘meaning attaches and interweaves itself with’ a collection of words. In general, the writer’s intention is to make the text as clear as possible, which in our thermodynamic formalism translates into a ‘pure state’, with von Neumann entropy equal to zero. Whether this pure state is also pure for the reader and/or listener of the text is an interesting question in itself about which, however, we will not elaborate in this article [47].
@@
Hence, in general, meaning, as used by people in human language, attaches and interweaves with states that are close to this Bose–Einstein condensate but usually not equal to it. It is an intriguing, but not yet experimentally solved question, whether things happen similarly in the micro-world in relation to ‘quantum coherence’, which would play an equivalent role to the notion of meaning in the human cognitive domain [47]. In any case, we believe that the analysis we have put forward in the present article comes closer to an explanation than the explanation produced by the symmetrization postulate and the ensuing complete indistinguishability of identical quantum physical entities.
@@
To conclude, we have observed that the classical statistical interpretation of thermodynamics rests on the assumption that physical entities behave in a random and independent way at a microscopic level, which entails maximization of entropy [1]. Based on our analysis of human language, we have suggested that it is more appropriate instead to consider that, also in the case of physics, there is no situation where entities are truly random and independent, as (conceptual or physical) entities always bundle (entangle) together, as evidenced by the presence of the Bose–Einstein statistics, with the aim to collaborate and reduce the overall uncertainty, hence entropy, and produce globally a pure state [42, 43]. Let us use this insight of the collaborative dynamics of words forming a story to complete our critique of the modern textbook introduction to quantum statistics. Indeed, we believe that this collaborative dynamics, that is, that words in a story work together and form their individual state accordingly, so that the whole story would be optimal in terms of the meaning the human mind is able to grasp in it, most fully defines the nature of words as quanta of human language. Thus, what follows is to be seen as a complement to what we have already presented in the previous pages, and in the articles cited above, considering in particular the issue of lack of independence, which has been of such burning concern in the early decades of QM.
Consider, e.g., the sentence Eleven Horses. We have already noted that in this sentence each of the horses is totally indistinguishable from the others. But is this statement really what is most important to characterize what is going on with each of the eleven horses in this sentence? It is much more relevant to note that Horse is an abstraction that refers to concrete horses, which can be of different types, with bodies located in space. But, still more relevant, is perhaps that Horse is a concept and not an object. Even more, suppose the above sentence is part of a somewhat longer sentence, namely the following: We Encountered Eleven Horses On Our Walk. Then, we see that the piece of sentence Eleven Horses changes by a mechanism we have called ‘contextual updating’, and now suddenly deals with horses in the neighbourhood where we were walking. But the mechanism of contextual updating is but one of the fantastically complex tools of subtle changes which words, the quanta of human language, are capable of. The most comprehensive way to grasp this complex adaptive dynamics of words is to state that ‘words collaborate and are willing to become more uncertain (more abstract), with the intention of making the story they are part of as certain (concrete) as possible’. And it is this intention, to minimize the entropy of the story the words are part of, that makes them choose the state in which they offer themselves. That a story consists of words some of which are identical, although they probably take on different roles in how they contribute to lowering the entropy of the whole story, considering also where in the story they appear, makes it possible for these identical words to be interchanged without changing the story, which is the ‘exchange symmetry’ present in human language.
We can describe, however, even more in depth the way in which words collaborate to build a sentence or a story, taking into account that in human language such collaboration aims to come as close as possible to carrying the meaning that the person wishes to communicate with that specific story. The mind of the person uttering it attempts to choose the words that do so in the most appropriate way. Would it make sense to try to grasp this subtle dynamics by assuming that words are completely indistinguishable? Clearly not. In the sentence Eleven Horses, it can be meaningfully asserted that the horses are indistinguishable, but the reason is the “sacrifice” that this sentence makes to the story in which it appears, with the intention of making that story what it is desired to be by the person uttering it. This happens because of the “force” exerted by the meaning to which the collaborative dynamic of words submits. In a structurally very similar way, a collaborative dynamics of molecules or photons submits to the “force” exerted by the quantum coherence of the domain they are part of.
Concerning in particular the choice of assigning equal probability to each micro-state, which is an aspect usually not explained in the standard formulation of modern textbooks, we believe that it is related to the contextual updating mechanism that submits each micro-state to the overall meaning of the story, or to the entire quantum coherence domain of the gas in question. We do not rule out the possibility that this choice is an expression of a universal averaging over different probabilities of occurrence, for each event. A refinement of experiments, in the future, might shed light on these subtle dynamics of molecules, or photons, and highlight some of the mechanisms suggested above, which we know are present in the case of human language.
@@
Our results clearly indicate that a theoretical direction towards the development of a quantum-type thermodynamics that is alternative to existing formal approaches to ‘quantum thermodynamics for physics’, none of which is satisfying [78, 79, 80], is a valuable aim to pursue. A possible direction for such an alternative theoretical approach rests on the notions of energy and entropy as defined in cognitive-linguistics domains, and we refer to [43] for more details about the development of a thermodynamics of human cognition and human culture.